# Sign-Problem-Free Nuclear Quantum Monte Carlo

### v1

@@ Acknowledgement
Charge density distributions (Fig. 11) and corresponding charge radii (Table 2) for the six target nuclei were calculated using the pinhole algorithm Elhatisari2017* . This method induces sign problems and is typically restricted to light nuclei ($A\leq 40$). Calculations incorporated the proton size $R_{p}=0.84\,\mathrm{fm}$ via convolution.
@@
We compare three published parameterizations: the FRDM Moller2016* (mac-mic), PC-PK1 Zhang2022* (RMF), and UNEDF1 PRC85-024304* (Skyrme). Standard deviations from experimental masses are computed for even-even nuclei with $N,Z\leq 28$ (Table 3). Phenomenological models show $\sim 2$-$3$ MeV RMS deviations in this region, improving significantly for heavier nuclei where mean-field contributions dominate. LAT-OPT1 achieves comparable precision ($\sim 3$ MeV RMS) despite fewer parameters, while mac-mic models yield smaller deviations with more adjustable parameters.
Table 3 also includes results from the Wigner-SU(4) interaction in Ref. PLB797-134863* , demonstrating significant improvement through: (1) systematic finite-volume effect elimination via larger boxes; (2) comprehensive imaginary-time extrapolation replacing fixed-$\tau$ calculations. The inclusion of spin-orbit coupling and efficient optimization further reduced deviations from $\sim 10$ MeV to $3$ MeV. Future extensions to larger training sets may further improve accuracy.
Complete numerical results (Table LABEL:tab:gse) compare LAT-OPT1 with mean-field calculations, listing Coulomb contributions separately. We also present the original data as separate figures for each nuclei. We show the results without the Coulomb energy, which can be restored by looking up Table LABEL:tab:gse.
@@
Table 3: Standard deviations $\sigma$ of ground-state binding energies for even-even nuclei ($N,Z\leq 28$) across mass models. Key: "Parameters" = adjustable parameter count; "Nuclei" = nuclei count for $\sigma$; $\sigma_{\rm all}$ = standard deviation over full nuclide chart.
@@
Macroscopic-microscopic FRDM Moller2016* $>30$ 1.15 69 0.56

### v2

@@ Acknowledgement
Sign problem for odd-even and odd-odd nuclei
We have examined the performance of LAT-OPT1 for various odd-even and odd-odd heavy nuclei. Table 2 presents the average sign $\langle e^{i\theta}\rangle$ for representative nuclei computed with $L_{t}=350$ using identical computational settings as for even-even nuclei. The selected nuclei include three odd-even systems (41Ca, 57Ni, 85Zr) and one odd-odd nucleus (82Nb). For all odd-even nuclei, the average sign exceeds $0.9$, indicating that the QMC statistical errors are comparable to those of neighboring even-even nuclei. For $^{82}\mathrm{Nb}$, where $\langle e^{i\theta}\rangle\approx 0.75$, the statistical uncertainty is merely $\sim 0.1\,\mathrm{MeV}$ larger than in adjacent even-even nuclei. We therefore conclude that our method enables precise extraction of ground-state energies for all nuclei with atomic mass numbers $A\leq 100$.
Table 2: Average sign $\langle e^{i\theta}\rangle$ for selected odd-even and odd-odd nuclei.
41Ca 57Ni 82Nb 85Zr
$\langle e^{i\theta}\rangle$ 0.915 0.920 0.750 0.927
@@
Charge density distributions (Fig. 12) and corresponding charge radii (Tab. 3) for the six target nuclei were calculated using the pinhole algorithm Elhatisari2017* . This method induces sign problems and is typically restricted to light nuclei ($A\leq 40$). Calculations incorporated the proton size $R_{p}=0.84\,\mathrm{fm}$ via convolution.
@@
We compare three published parameterizations: the FRDM Moller2016* (mac-mic), PC-PK1 Zhang2022* (RMF), and UNEDF1 PRC85-024304* (Skyrme). Standard deviations from experimental masses are computed for even-even nuclei with $N,Z\leq 28$ (Tab. 4). Phenomenological models show $\sim 2$-33 MeV RMS deviations in this region, improving significantly for heavier nuclei where mean-field contributions dominate. LAT-OPT1 achieves comparable precision ($\sim 3$ MeV RMS) despite fewer parameters, while mac-mic models yield smaller deviations with more adjustable parameters.
Tab. 4 also includes results from the Wigner-SU(4) interaction in Ref. PLB797-134863* , demonstrating significant improvement through: (1) systematic finite-volume effect elimination via larger boxes; (2) comprehensive imaginary-time extrapolation replacing fixed-$\tau$ calculations. The inclusion of spin-orbit coupling and efficient optimization further reduced deviations from $\sim 10$ MeV to $3$ MeV. Future extensions to larger training sets may further improve accuracy.
Complete numerical results (Tab. LABEL:tab:gse) compare LAT-OPT1 with mean-field calculations, listing Coulomb contributions separately. We also present the original data as separate figures for each nuclei. We show the results without the Coulomb energy, which can be restored by looking up Tab. LABEL:tab:gse.
@@
Table 4: Standard deviations $\sigma$ of ground-state binding energies for even-even nuclei ($N,Z\leq 28$) across mass models. Here, Parameters is adjustable parameter count, Nuclei is nuclei count for $\sigma$, and $\sigma_{\rm all}$ is standard deviation over full nuclide chart.
@@
Macroscopic-microscopic FRDM Moller2016* >30>30 1.15 69 0.56